OpenStudy (anonymous):

please help me factor: x^2+2xy+2x+y^2+2y-8 How come we dont factor out 8 with 2?

OpenStudy (freckles):

big hint: (x+y)^2 =?

OpenStudy (anonymous):

I know but the other part I got wrong

OpenStudy (freckles):

what is (x+y)^2=?

OpenStudy (freckles):

and remember (x+y)^2=(x+y)(x+y)=

OpenStudy (freckles):

when you multiply that out what do you have

OpenStudy (anonymous):

(x+y)^2 +2(x+y) - 8 => x+y = a a^2 + 2a - 8 = a^2 +4a -2a -8 = (a+4)(a-2) = (x+y+4)(x+y-2)

OpenStudy (anonymous):

Hope its Correct...